This is a test for BibTeX Entry of my publications. [1] [2] [3] [4] [5] [6] [7]

[7] [8] [9] [10] [11] [12] [13] [14] [15] [16] [17]

References

[1] M. Amintoosi, F. Farbiz, and M. Fathy, “A QR Decomposition based mixture model algorithm for background modeling,” in ICICS2007, Sixth International Conference on Information, Communication and Signal Processing, Singapore, December 2007, pp. 1–5.

[2] M. Amintoosi, H. S. Yazdi, and J. Haddadnnia, “Fuzzy student sectioning,” in PATAT04: Practice and Theory of Automated Timetabling, USA, Aug 2004, pp. 421–424.

[3] M. Amintoosi and H. S. Yazdi, “Optimum fuzzy student sectioning with a fuzzy evaluation function in a genetic course timetabling,” in Proceedings of the 9th Iranian Annual International CSI Computer Conference, Sharif Univ. of Tech.,Iran, Feb 2004, pp. 345–352, persian.

[4] M. Amintoosi, R. Monsefi, and J. Haddadnia, “Using pattern matching for tiling and packing problems,” in Fifth International Conference on Computer Sciences, ser. Modeling, Computation and Optimization in Information Systems and Management Sciences. Metz,France: Hermes Science Publishing, July 2004, pp. 97–104.

[5] M. Amintoosi, M. Fathy, N. Mozayani, and A. Rahmani, “A fish school clustering algorithm: Applied to student sectioning problem,” Dynamics of Continuous Discrete & Impulse Systems, series B: Applications and Algorithms, vol. 2, pp. 696–699, December 2007, post Proceeding of LSMS2007, Life System Modeling and Simulation 2007, China.

[6] M. Amintoosi and J. Haddadnnia, “Feature selection in a fuzzy student sectioning algorithm,” in LNCS, ser. Lecture Notes in Computer Science. Springer- Verleg, 2005, vol. 3616, pp. 147–160.

[7] M. Amintoosi, F. Farbiz, M. Fathy, M. Analoui, and N. Mozayani, “QR-Decomposition-based Algorithm for Background Subtraction,” in ICASSP2007, IEEE International Conference on Acoustics, Speech, and Signal Processing, USA, 2007, pp. 1093–1096.

[8] M. Amintoosi, H. S. Yazdi, M.Fathy, and R. Monsefi, “Using pattern matching for tiling and packing problems,” European Journal of Operational Research, vol. 183, pp. 950–960, 2007.

[9] M. Amintoosi, N. Mozayani, and M. Fathy, “Spatially varying image superresolution,” in (MVIP08) 2008 International Conference on Machine Vision and Image Processing, Tabriz,Iran, 2008, submitted.

[10] A. Amiri, M. Fathy, M. Amintoosi, and H. S. Yazdi, “Modified quantized input variable step size LMS, QX-VSS LMS algorithm applied to signal prediction,” in Proceeidings of 4th IEEE GCC Conference, -, November 2007.

[11] A. Amiri, M. Fathy, M. Amintoosi, and H. Sadoghi, “A new quantized input RLS, QI-RLS, algorithm,” in ICCSA (3), 2007, pp. 495–506.

[12] M. Fathy, M. Analoui, N. Mozayani, M. Amintoosi, and F. Farbiz, “A background model initialization algorithm based on qr-decomposition,” in 4th Iranian Conference on Machine Vision and Image Processing, Mashhad,Iran, Feb 2007.

[13] M. Fathy, N. Mozayani, and M. Amintoosi, “Outlier removal for super-resolution problem using QR-Decomposition,” in 2008 International Conference on Image Processing, Computer Vision, and Pattern Recognition (IPCV08), USA, July 2008, accepted.

[14] H. S. Yazdi, M. Amintoosi, and M.Fathy, “A new spatio-temporal motion stacking and using it for facial expression recognition,” in The First Local Conference on Electrical Engineering of Islamic Azad University, Gonabad,Iran, 2006, persian.

[15] H. S. Yazdi, M. Amintoosi, and M. Fathy, “Facial expression recognition with qim and itmi spatio-temporal database,” in 4th Iranian Conference on Machine Vision and Image Processing, Mashhad,Iran, Feb 2007, pp. –.

[16] R. Monsefi and M. Amintoosi, “A genetic-neuro algorithm for tiling problems,” in Proceedings of 5th Iranian Annual International CSI Computer Conference, Shahid Beheshti Univ.,Iran, March 2000, pp. 298–304, persian.

[17] ——, “A genetic-neuro algorithm for tiling problems with rotation and reflection of figures,” Iranian Journal of Science and Technology, vol. 26, no. B4, pp. 693– 700, December 2002.